# Bitcoin Under Volatile Block Rewards: How Mempool Statistics Can Influence Bitcoin Mining

### v2

@@ Appendix C Non-Predictable WeRLman
1:Input: Previous state element prev_state.
2:Output: Updated states new_with_fee, new_without_fee.
3:$current\_max\_fork\leftarrow\max(h,a)$
@@
5:Create a copy of the current state: $new\_with\_fee\leftarrow\text{copy}()$
6:Create another copy of the current state: $new\_without\_fee\leftarrow\text{copy}()$
@@
8:     Increment the pool of new_with_fee: $new\_with\_fee.pool\leftarrow new\_with\_fee.pool+1$
[…]
10:     if $new\_honest$ then
[…]
12:         Increment T_h of new_with_fee: $new\_with\_fee.T\_h\leftarrow new\_with\_fee.T\_h+with\_fee$
@@
[…]
16:              Increment T_a of new_with_fee: $new\_with\_fee.T\_a\leftarrow new\_with\_fee.T\_a+with\_fee$
@@
18:              Set the last element of L: $new\_with\_fee.L[-1]\leftarrow 1$
@@
22:Return new_with_fee, new_without_fee
Appendix D Adversarial Strategies in the Original and Non-predictable WeRLman Environments
@@
Let $n_{\mathcal{A}}$, $n_{\mathcal{H}}$ denote the number of blocks in the adversarial and the honest fork, respectively. Let $\alpha$, $g$, $P$, and $F$ denote the adversarial mining share, the adversarial communication capability, the whale transaction probability, and the extra fee of a whale transaction, respectively. We denote by $N_{\mathcal{A}}$, $N_{\mathcal{H}}$, and $R_{\mathcal{A}}$ the normalized average number of canonical adversarial blocks, the normalized average number of canonical honest blocks, and the normalized average adversarial block reward. The time-averaged profit of strategy $\pi$ can be calculated as $\texttt{Profit}(\pi)=\frac{R_{\mathcal{A}}{(\pi)}}{N_{\mathcal{A}}{(\pi)}+N_{\mathcal{H}}{(\pi)}}$. Note that the time-averaged profit of the adversary if mining honestly is equal to $\texttt{Profit}(\pi^{\mathcal{H}})=\alpha(1_{F}P)$.
D.1. Strategy $\pi_{1}^{\text{WeRLman}}$
Let $\texttt{tnx}=0$ denote the existence of a normal transaction in the pool, and $\texttt{tnx}=1$ denote the existence of a whale transaction in the pool. The state space and the action in each state are as follows:
State $S_{0}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0,\texttt{tnx}=0)$: If the next block is honest: adopt. If the next block is adversarial and the next transaction is a whale: wait. If the next block is adversarial and the next transaction is normal: publish.
State $S_{0}^{\prime}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0,\texttt{tnx}=1)$: If the next block is honest: adopt. If the next block is adversarial: publish.
State $S_{1}:(n_{\mathcal{A}}=1(\texttt{normal}),n_{\mathcal{H}}=0,\texttt{tnx}=1)$: If the next block is adversarial: wait until $n_{\mathcal{A}}-n_{\mathcal{H}}=1$, then Override. If the next block is honest: match.
@@ D.1. Strategy $\pi_{1}^{\text{WeRLman}}$
State $S_{2}:(n_{\mathcal{A}}=1(\texttt{normal-published}),n_{\mathcal{H}}=1)$: In this state, a new transaction is not sampled, as the same height block already exists. If the next block is honest: adopt. If the next block is adversarial: publish.
@@
Let $P_{0}$, $P_{0}^{\prime}$, $P_{1}$, $P_{2}$ denote the stationary probability of being in states $S_{0}$, $S_{0}^{\prime}$, $S_{1}$, $S_{2}$, respectively. We have:
(11) $\begin{split}&P_{0}=\frac{1-p}{1+p(1-p)\alpha(1-\alpha)},\quad P_{1}=\frac{\alpha p(1-p)}{1+p(1-p)\alpha(1-\alpha)},\quad\\ &P_{2}=\frac{p(1-p)\alpha(1-\alpha)}{1+p(1-p)\alpha(1-\alpha)},\quad P_{0}^{\prime}=1-P_{0}-P_{1}-P_{2}.\end{split}$
@@
The normalized average number of canonical honest blocks, canonical adversarial blocks, can be obtained as follows:
@@
(12) $\begin{split}N_{\mathcal{H}}=&P_{0}\left(1-\alpha\right)+2P_{2}\left(1-\alpha\right)\left(1-g\right)+P_{2}\left(1-\alpha\right)g+P_{0}^{\prime}\left(1-\alpha\right),\\ N_{\mathcal{A}}=&P_{0}\alpha\left(1-p\right)+\left(2+\frac{\alpha}{1-2\alpha}\right)P_{1}\alpha+2P_{2}\alpha+P_{2}\left(1-\alpha\right)g+P_{0}^{\prime}\alpha,\\ R_{\mathcal{A}}=&P_{0}\alpha\left(1-p\right)+P_{0}^{\prime}\alpha\left(1+F\right)+P_{2}\alpha\left(2+F\right)+P_{2}\left(1-\alpha\right)g\\ &+P_{1}\alpha\left(2+F\right)+P_{1}\alpha\frac{\alpha}{1-2\alpha}\left(1+pF\right)\enspace.\end{split}$
D.2. Strategy $\pi_{1}^{\text{Non-predictable}}$
Let $\texttt{tnx}=0$ denote if the transaction included in the adversarial block is a normal one. The state space and the action in each state are as follows:
State $S_{0}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0)$: If the next block is honest: adopt. If the next block is adversarial and includes a whale: publish. If the next block is adversarial and includes a normal transaction: wait.
State $S_{1}:(n_{\mathcal{A}}=1(\texttt{normal}),n_{\mathcal{H}}=0,\texttt{tnx}=0)$: If the next block is honest: match. If the next block is adversarial: publish.
State $S_{2}:(n_{\mathcal{A}}=1(\texttt{normal-published}),n_{\mathcal{H}}=1,\texttt{tnx}=0)$: If the next block is honest: adopt. If the next block is adversarial: publish.
Let $P_{0}$, $P_{1}$, $P_{2}$ denote the stationary probability of being in states $S_{0}$, $S_{1}$, and $S_{2}$, respectively. We have:
(13) $\begin{split}&P_{0}=\frac{1}{1+\left(1-p\right)\alpha\left(2-\alpha\right)},\quad P_{1}=\alpha\left(1-p\right)P_{0},\quad\\ &P_{2}=\left(1-\alpha\right)P_{1}.\end{split}$
@@ D.2. Strategy $\pi_{1}^{\text{Non-predictable}}$
The normalized average number of canonical honest blocks, canonical adversarial blocks, can be obtained as follows:
@@
(14) $\begin{split}N_{\mathcal{H}}=&P_{0}\left(1-\alpha\right)+P_{2}\left(1-\alpha\right)g+2P_{2}\left(1-\alpha\right)\left(1-g\right),\\ N_{\mathcal{A}}=&P_{0}\alpha p+P_{1}\alpha\left(2+\frac{\alpha}{1-2\alpha}\right)+2P_{2}\alpha+P_{2}\left(1-\alpha\right)g,\\ R_{\mathcal{A}}=&P_{0}\alpha p\left(1+F\right)+P_{1}\alpha\left(1+\left(1+pF\right)\left(1+\frac{\alpha}{1-2\alpha}\right)\right)\\ &+P_{2}\left(1-\alpha\right)g+P_{2}\alpha\left(2+pF\right)\enspace.\end{split}$
D.3. Strategy $\pi_{2}^{\text{Non-predictable}}$
Let $\texttt{tnx}=1$ denote if the transaction included in the honest block is a whale transaction. The state space and the action in each state are as follows:
State $S_{0}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0)$: If the next block is honest and includes a normal transaction: adopt. If the next block is honest and includes a whale transaction: wait. If the next block is adversarial: publish.
State $S_{1}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=1,\texttt{tnx}=1)$: If the next block is honest and includes a normal transaction: adopt. If the next block is honest and includes a whale transaction: adopt the first honest block. If the next block is adversarial: wait.
State $S_{2}:(n_{\mathcal{A}}=1(\texttt{whale-unpublished}),n_{\mathcal{H}}=1)$: If the next block is honest and includes a normal transaction: adopt. If the next block is honest and includes a whale transaction: adopt the first honest block. If the next block is adversarial: publish.

### v3

@@ Appendix B Non-Predictable WeRLman
1:Input: Current state CSCS (self), and previous state $PS$.
2:Output: Updated states new_with_feenew\_with\_fee, new_without_feenew\_without\_fee.
@@
5:Create a copy of the current state: $new\_with\_fee\leftarrow\text{copy}(CS)$
6:Create another copy of the current state: $new\_without\_fee\leftarrow\text{copy}(CS)$
7:if prev_max_fork<current_max_forkprev\_max\_fork<current\_max\_fork and CS.pool<max_poolCS.pool<max\_pool then
8:  $new\_with\_fee.pool\leftarrow new\_with\_fee.pool+1$
9:  $new\_honest\leftarrow\text{True}$ if (CS.hPS.h)>0(CS.h-PS.h)>0 else False
10:  if new_honestnew\_honest then
11:   $with\_fee\leftarrow 1$ if CS.T_h=PS.T_hCS.T\_h=PS.T\_h else 0
12:   $new\_with\_fee.T\_h\leftarrow new\_with\_fee.T\_h+with\_fee$
@@
15:     $with\_fee\leftarrow 1$ if CS.T_a=PS.T_aCS.T\_a=PS.T\_a else 0
16:     $new\_with\_fee.T\_a\leftarrow new\_with\_fee.T\_a+with\_fee$
@@
18:     $new\_with\_fee.L[-1]\leftarrow 1$
@@
22:Return new_with_feenew\_with\_fee, new_without_feenew\_without\_fee
Appendix C Adversarial Strategies in the Original and Non-predictable WeRLman Environments
@@
Let $\alpha$, $g$, $p$, and $F$ denote the adversarial mining share, the adversarial communication capability, the whale transaction probability, and the extra fee of a whale transaction, respectively. We denote by $_{\mathcal{A}}, NN_{\mathcal{H}}$, and $R_{\mathcal{A}}$ the normalized average number of canonical adversarial blocks, the normalized average number of canonical honest blocks, and the normalized average adversarial block reward. The time-averaged profit of strategy $\pi$ can be calculated as $\texttt{Profit}(\pi)=\frac{R_{\mathcal{A}}{(\pi)}}{N_{\mathcal{A}}{(\pi)}+N_{\mathcal{H}}{(\pi)}}$. Note that the time-averaged profit of the adversary if mining honestly is equal to $\texttt{Profit}(\pi^{\mathcal{H}})=\alpha(1+Fp)$.
We introduce three strategies: $\pi_{1}^{\text{WeRLman}}$, $\pi_{1}^{\text{Non-predictable}}$, and $\pi_{2}^{\text{Non-predictable}}$. Each strategy specifies the adversary’s action in each state of the fork race. The adversary takes these actions upon the event of block mining. A newly mined block is adversarial with probability $\alpha$ and honest with probability $1-\alpha$. At each event of mining a block at a new height, a transaction is sampled: it is a whale transaction with probability $p$ and a normal transaction with probability 1p1-p. Note that in the original WeRLman environment (used by strategy $\pi_{1}^{\text{WeRLman}}$), the sampled transaction is added to the mempool and included in the next block. In contrast, in the non-predictable WeRLman environment (used by strategies $\pi_{1}^{\text{Non-predictable}}$ and $\pi_{2}^{\text{Non-predictable}}$), the sampled transaction is assumed to have already been included in the newly mined block. In a same-height fork race between a published adversarial fork and an honest fork, in addition to the adversarial mining power, a fraction $g$ of honest miners will also mine on top of the adversarial fork.
We use four actions to describe the strategies: adopt, denoting acceptance of the honest fork as the canonical chain; wait, denoting continued mining on top of the adversarial fork; publish, denoting the release of a longer adversarial fork; and match, denoting the release of an adversarial fork at the same height as the honest fork immediately after the publication of a new honest block.
C.1. Strategy $\pi_{1}^{\text{WeRLman}}$
We use a tuple $(n_{\mathcal{A}},n_{\mathcal{H}},\texttt{whale}^{\texttt{pool}},\texttt{whale}^{\texttt{block}})$ to represent a state during a fork race. Here, $n_{\mathcal{A}}$ and $n_{\mathcal{H}}$ denote the number of blocks in the adversarial and honest forks, respectively. The variable $\texttt{whale}^{\texttt{pool}}$ indicates the type of transaction currently in the pool w.r.t. the adversarial fork, which can be included in the next block: $\texttt{whale}^{\texttt{pool}}=1$ if there is a whale transaction, and $\texttt{whale}^{\texttt{pool}}=0$ if there is a normal transaction. $\texttt{whale}^{\texttt{block}}$ indicates the type of transaction included in the adversarial block: $\texttt{whale}^{\texttt{block}}=1$ if the block includes a whale transaction, and $\texttt{whale}^{\texttt{block}}=0$ if it includes a normal transaction.
We define four states as follows:
  • State $S_{0,0}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0,\texttt{whale}^{\texttt{pool}}=0,\texttt{whale}^{\texttt{block}}=\bot)$. This state denotes that both the adversary and honest miners are mining on top of the same canonical chain, and the existing transaction in the mempool is normal.
  • State $S_{0,0}^{\prime}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0,\texttt{whale}^{\texttt{pool}}=1,\texttt{whale}^{\texttt{block}}=\bot)$. This state denotes that both the adversary and honest miners are mining on top of the same canonical chain, but the existing transaction in the mempool is a whale transaction.
  • State $S_{1,0}:(n_{\mathcal{A}}=1,n_{\mathcal{H}}=0,\texttt{whale}^{\texttt{pool}}=1,\texttt{whale}^{\texttt{block}}=0)$. This state denotes that the secret adversarial fork contains a single block with a normal transaction, while a whale transaction remains in the mempool to be included in the next block.
@@ C.1. Strategy $\pi_{1}^{\text{WeRLman}}$
  • State $S_{1,1}:(n_{\mathcal{A}}=1,n_{\mathcal{H}}=1,\texttt{whale}^{\texttt{pool}}=1,\texttt{whale}^{\texttt{block}}=0)$. This state denotes that there is a fork race between the adversarial and honest forks, both with a height of one block. The adversarial block includes a normal transaction, while the honest block includes a whale transaction. The next block in the adversarial fork can steal the whale transaction from the honest block. Note that upon transferring to this state, a new transaction is not sampled, as the same height block already exists.
Strategy $\pi_{1}^{\text{WeRLman}}$ is introduced in Table 1.
Table 1. Strategy $\pi_{1}^{\text{WeRLman}}$.
Initial State Action Probability Final State Reward
$S_{0,0}$
wait.
If the next block is honest, then switch to adopt.
If the next block is adversarial and the next transaction
is normal, then switch to publish.
$\alpha p$ $S_{1,0}$ 0
$\alpha(1-p)$ $S_{0,0}$ 1
$(1-\alpha)p$ $S_{0,0}^{\prime}$ 0
$(1-\alpha)(1-p)$ $S_{0,0}$ 0
$S_{0,0}^{\prime}$
wait.
If next block is honest, then switch to adopt.
If the next block is adversarial, then switch to publish.
$\alpha p$ $S_{0,0}^{\prime}$ $1+F$
$\alpha(1-p)$ $S_{0,0}$ $1+F$
$(1-\alpha)p$ $S_{0,0}^{\prime}$ 0
$(1-\alpha)(1-p)$ $S_{0,0}$ 0
$S_{1,0}$
wait.
If the next block is adversarial, wait until $n_{\mathcal{A}}-n_{\mathcal{H}}=1$,
then switch to publish.
$\alpha p$ $S_{0,0}^{\prime}$ $2+F+(1+pF)\frac{\alpha}{1-2\alpha}$
$\alpha(1-p)$ $S_{0,0}^{\prime}$ $2+F+(1+pF)\frac{\alpha}{1-2\alpha}$
$(1-\alpha)$ $S_{1,1}$ 0
$S_{1,1}$ match $\alpha p$ $S_{0,0}^{\prime}$ $2+F$
$\alpha(1-p)$ $S_{0,0}$ $2+F$
$(1-\alpha)gp$ $S_{0,0}^{\prime}$ 1
$(1-\alpha)g(1-p)$ $S_{0,0}$ 1
$(1-\alpha)(1-g)p$ $S_{0,0}^{\prime}$ 0
$(1-\alpha)(1-g)(1-p)$ $S_{0,0}$ 0
@@
Let $P_{0,0}$, $P_{0,0}^{\prime}$, $P_{1,0}$, $P_{1,1}$ denote the stationary probability of being in states $S_{0,0}$, $S_{0,0}^{\prime}$, $S_{1,0}$, $S_{1,1}$, respectively. We have:
(11) $\begin{split}&P_{0,0}=\frac{1-p}{1+p(1-p)\alpha(1-\alpha)},\quad P_{1,0}=\frac{\alpha p(1-p)}{1+p(1-p)\alpha(1-\alpha)},\quad\\ &P_{1,1}=\frac{p(1-p)\alpha(1-\alpha)}{1+p(1-p)\alpha(1-\alpha)},\quad P_{0,0}^{\prime}=1-P_{0,0}-P_{1,0}-P_{1,1}.\end{split}$
@@
The normalized numbers of canonical honest blocks, canonical adversarial blocks, and the normalized adversarial block reward achieved under strategy $\pi_{1}^{\text{WeRLman}}$ can be obtained as follows:
@@
(12) $\begin{split}N_{\mathcal{H}}=&P_{0,0}\left(1-\alpha\right)+2P_{1,1}\left(1-\alpha\right)\left(1-g\right)\\ &+P_{1,1}\left(1-\alpha\right)g+P_{0,0}^{\prime}\left(1-\alpha\right),\\ N_{\mathcal{A}}=&P_{0,0}\alpha\left(1-p\right)+\left(2+\frac{\alpha}{1-2\alpha}\right)P_{1,0}\alpha+2P_{1,1}\alpha\\ &+P_{1,1}\left(1-\alpha\right)g+P_{0,0}^{\prime}\alpha,\\ R_{\mathcal{A}}=&P_{0,0}\alpha\left(1-p\right)+P_{0,0}^{\prime}\alpha\left(1+F\right)+P_{1,1}\alpha\left(2+F\right)+P_{1,1}\left(1-\alpha\right)g\\ &+P_{1,0}\alpha\left(2+F\right)+P_{1,0}\alpha\frac{\alpha}{1-2\alpha}\left(1+pF\right)\enspace.\end{split}$
C.2. Strategy $\pi_{1}^{\text{Non-predictable}}$
We represent a state during a fork race using the following tuple: $(n_{\mathcal{A}},n_{\mathcal{H}},\texttt{whale}^{\texttt{block}})$. The descriptions of these parameters are the same as those provided in Appendix C.1. We define three states as follows:
  • State $S_{0,0}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0,\texttt{whale}^{\texttt{block}}=\bot)$. This state denotes that both the adversary and honest miners are mining on top of the same canonical chain.
  • State $S_{1,0}:(n_{\mathcal{A}}=1,n_{\mathcal{H}}=0,\texttt{whale}^{\texttt{block}}=0)$. This state denotes that the secret adversarial fork contains a single block with a normal transaction.
  • State $S_{1,1}:(n_{\mathcal{A}}=1,n_{\mathcal{H}}=1,\texttt{whale}^{\texttt{block}}=0)$. This state denotes that there is a fork race between the adversarial and honest forks, both with a height of one block. The adversarial block includes a normal transaction. Note that upon transferring to this state, a new transaction is not sampled, as the same height block already exists.
Strategy $\pi_{1}^{\text{Non-predictable}}$ is presented in Table 2.
Table 2. Strategy $\pi_{1}^{\text{Non-predictable}}$.
Initial State Action Probability Final State Reward
$S_{0,0}$
wait.
If the next block is honest, then switch to adopt.
If the next block is adversarial with a whale, then publish.
$\alpha p$ $S_{0,0}$ $1+F$
$\alpha(1-p)$ $S_{1,0}$ 0
$(1-\alpha)$ $S_{0,0}$ 0
$S_{1,0}$
wait.
If the next block is adversarial, wait until $n_{\mathcal{A}}-n_{\mathcal{H}}=1$,
then switch to publish.
$\alpha$ $S_{0,0}$ $1+(1+pF)\left(1+\frac{\alpha}{1-2\alpha}\right)$
$(1-\alpha)$ $S_{1,1}$ 0
$S_{1,1}$ match $\alpha p$ $S_{0,0}$ $2+F$
$\alpha(1-p)$ $S_{0,0}$ 22
$(1-\alpha)g$ $S_{0,0}$ 1
$(1-\alpha)(1-g)$ $S_{0,0}$ 0
Let $P_{0,0}$, $P_{1,0}$, $P_{1,1}$ denote the stationary probability of being in states $S_{0,0}$, $S_{1,0}$, and $S_{1,1}$, respectively. We have:
(13) $\begin{split}&P_{0,0}=\frac{1}{1+\left(1-p\right)\alpha\left(2-\alpha\right)},\quad P_{1,0}=\alpha\left(1-p\right)P_{0,0},\quad\\ &P_{1,1}=\left(1-\alpha\right)P_{1,0}.\end{split}$
@@ C.2. Strategy $\pi_{1}^{\text{Non-predictable}}$
The normalized numbers of canonical honest blocks, canonical adversarial blocks, and the normalized adversarial block reward achieved under strategy $\pi_{1}^{\text{Non-predictable}}$ can be obtained as follows:
@@
(14) $\begin{split}N_{\mathcal{H}}=&P_{0,0}\left(1-\alpha\right)+P_{1,1}\left(1-\alpha\right)g+2P_{1,1}\left(1-\alpha\right)\left(1-g\right),\\ N_{\mathcal{A}}=&P_{0,0}\alpha p+P_{1,0}\alpha\left(2+\frac{\alpha}{1-2\alpha}\right)+2P_{1,1}\alpha+P_{1,1}\left(1-\alpha\right)g,\\ R_{\mathcal{A}}=&P_{0,0}\alpha p\left(1+F\right)+P_{1,0}\alpha\left(1+\left(1+pF\right)\left(1+\frac{\alpha}{1-2\alpha}\right)\right)\\ &+P_{1,1}\left(1-\alpha\right)g+P_{1,1}\alpha\left(2+pF\right)\enspace.\end{split}$
C.3. Strategy $\pi_{2}^{\text{Non-predictable}}$
We represent a state during a fork race using the following tuple: $(n_{\mathcal{A}},n_{\mathcal{H}},\texttt{whale}^{\texttt{block}}_{\mathcal{H}})$. Here, $n_{\mathcal{A}}$ and $n_{\mathcal{H}}$ denote the number of blocks in the adversarial and honest forks, respectively. $\texttt{whale}^{\texttt{block}}_{\mathcal{H}}$ indicates the type of transaction included in the first block of the honest fork: $\texttt{whale}^{\texttt{block}}_{\mathcal{H}}=1$ if the block includes a whale transaction, and $\texttt{whale}^{\texttt{block}}_{\mathcal{H}}=0$ if it includes a normal transaction. We define the states as follows:
  • State $S_{0,0}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0,\texttt{whale}^{\texttt{block}}_{\mathcal{H}}=\bot)$. This state denotes that both the adversary and honest miners are mining on top of the same canonical chain.
  • State $S_{0,1}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=1,\texttt{whale}^{\texttt{block}}_{\mathcal{H}}=1)$. This state denotes that the honest fork contains a single block with a whale transaction.
  • State $S_{i,i}:(n_{\mathcal{A}}=i,n_{\mathcal{H}}=i,\texttt{whale}^{\texttt{block}}=1)$ for $i\geq 1$. This state denotes that there is a fork race between the adversarial and honest forks, both with a height of $i$ blocks. The first block in the honest fork includes a whale. The action match is not feasible in this state, as the $i^{\text{th}}$ adversarial block is mined after the $i^{\text{th}}$ honest block. Note that upon transferring to this state, a new transaction is not sampled, as the same height block already exists.
  • State $S_{i,i+1}:(n_{\mathcal{A}}=i,n_{\mathcal{H}}=i+1,\texttt{whale}^{\texttt{block}}=1)$ for $i\geq 1$. This state denotes that there is a fork race between the adversarial fork of height $i$ blocks with an honest fork of height $i+1$ blocks. The first block in the honest fork includes a whale.